Given the following demand equation, Qd = 120-6P, what it the quantity demanded when the price is 3?
Answer #1

As per the given quantity demand equation above :-

Qd = 120 - 6P

We need to find out quantity demanded when the price is 3. So puuting the value of price as P=3 , we get the value of quantity demanded as :-

Qd = 120 - 6(3)

Qd = 120 - 18

Qd = 102

So the quantity demanded is 102 when the price is 3.